Selected T-accounts of Moore Company are given below for the just completed year:

Raw Materials
Bal. 1/1 27,000 Credits ?
Debits 144,000
Bal. 12/31 37,000
Manufacturing Overhead
Debits 197,480 Credits ?
Work in Process
Bal. 1/1 32,000 Credits 506,000
Direct materials 102,000
Direct labor 186,000
Overhead 219,480
Bal. 12/31 ?
Factory Wages Payable
Debits 209,000 Bal. 1/1 15,000
Credits 204,000
Bal. 12/31 10,000
Finished Goods
Bal. 1/1 52,000 Credits ?
Debits ?
Bal. 12/31 78,000
Cost of Goods Sold
Debits ?

Q.

Compute the ending balance in Work in Process. Assume that this balance consists entirely of goods started during the year. If $11,000 of this balance is direct labor cost, how much of it is direct materials cost? Applied overhead cost?

Ending balance in the work in process

Direct Material Cost

Applied Overhead Cost
Answer #1

Ending balance in work in process = 32000+102000+186000+219480-506000 = 33480

Direct material = 33480-12980-11000 = 9500

Applied overhead = 11000*1.18 = 12980